1- Exercise 11-06 Crane Company purchased equipment for $192,000 on October 1, 2020. It is estimated...

1-

Exercise 11-06

Crane Company purchased equipment for $192,000 on October 1, 2020. It is estimated that the equipment will have a useful life of 8 years and a salvage value of $12,000. Estimated production is 36,000 units and estimated working hours are 20,000. During 2020, Crane uses the equipment for 500 hours and the equipment produces 1,000 units.

Compute depreciation expense under each of the following methods. Crane is on a calendar-year basis ending December 31. (Round rate per hour and rate per unit to 2 decimal places, e.g. 5.35 and final answers to 0 decimal places, e.g. 45,892.)

(a)

Straight-line method for 2020

$enter a dollar amount

(b)

Activity method (units of output) for 2020

$enter a dollar amount

(c)

Activity method (working hours) for 2020

$enter a dollar amount

(d)

Sum-of-the-years'-digits method for 2022

$enter a dollar amount

(e)

Double-declining-balance method for 2021

$enter a dollar amount
Answer #1
Please give positive ratings so I can keep answering. Thanks!
Crane Company
Cost of Equipment    192,000.00 A
Residual Value      12,000.00 B
Depreciable Value    180,000.00 C=A-B
Answer a
Straight Line Method
Depreciable Value    180,000.00 See C
Life                8.00 D
Annual depreciation      22,500.00 E=C/D
Purchased on Oct 1, 2020
Depreciation for 2020 (months)                3.00 F
Depreciation expense for 2020        5,625.00 G=E/12*F
Answer b
Units of output Method
Depreciable Value    180,000.00 See C
Life (Units)      36,000.00 H
Depreciation per unit                5.00 I=C/H
Production in 2020        1,000.00 J
Depreciation expense for 2020        5,000.00 K=I*J
Answer c
Working hours Method
Depreciable Value    180,000.00 See C
Life (hours)      20,000.00 L
Depreciation per hour                9.00 M=C/L
Hours used in 2020            500.00 N
Depreciation expense for 2020        4,500.00 O=M*N
C P Q=C*P
Answer d Table Depreciable Value Remaining life Depreciable fraction Depreciable Expense
Sum-of-the-years'-digits method Year 2020              180,000.00                8.00 8/36     40,000.00
Depreciable Value    180,000.00 See C Year 2021              180,000.00                7.00 7/36     35,000.00
Life (years)                8.00 See D Year 2022              180,000.00                6.00 6/36     30,000.00
Refer table in the right side. Year 2023              180,000.00                5.00 5/36     25,000.00
That is annual depreciation. Year 2024              180,000.00                4.00 4/36     20,000.00
But equioment purchased on Oct 1 Year 2025              180,000.00                3.00 3/36     15,000.00
Depreciation for 2020 (months)                3.00 See F Year 2026              180,000.00                2.00 2/36     10,000.00
Annual Depreciation for 2020      40,000.00 Z Year 2027              180,000.00                1.00 1/36       5,000.00
Depreciation expense for 2020      10,000.00 AA= Z/12*F              36.00 180,000.00
Remaining will be allicated to 2021      30,000.00 AB= Z-AA
Annual Depreciation for 2021      35,000.00 AC
3 months deprciation        8,750.00 AD= AC/12*F
Depreciation expense for 2021      38,750.00 AE= AB+AD
Remaining will be allicated to 2022      26,250.00 AF=AC-AD
Annual Depreciation for 2022      30,000.00 AG
3 months deprciation        7,500.00 AH=AG/12*3
Depreciation expense for 2022      33,750.00 AI= AF+AH
Answer e
Double Declining Method
Cost of Equipment    192,000.00 See A
Life                8.00 See D
Annual depreciation      24,000.00 R=A/D
Depreciation rate 12.50% S=R/A
Double Depreciation % 25.00% T=S*2
Annual depreciation for 2020      48,000.00 U=A*T
Purchased on Oct 1, 2020
Depreciation for 2020 (months)                3.00 V
Depreciation expense for 2020      12,000.00 W=U/12*V
Book value in 2021    180,000.00 X=A-W
Depreciation expense for 2021      45,000.00 Y=X*T